# DiffusionNAG: Predictor-guided Neural Architecture Generation with Diffusion Models

### 31 Dec 2023 v2

@@ Abstract
Existing NAS methods suffer from either an excessive amount of time for repetitive sampling and training of many task-irrelevant architectures. To tackle such limitations of existing NAS methods, we propose a paradigm shift from NAS to a novel conditional Neural Architecture Generation (NAG) framework based on diffusion models, dubbed DiffusionNAG. Specifically, we consider the neural architectures as directed graphs and propose a graph diffusion model for generating them. Moreover, with the guidance of parameterized predictors, DiffusionNAG can flexibly generate task-optimal architectures with the desired properties for diverse tasks, by sampling from a region that is more likely to satisfy the properties. This conditional NAG scheme is significantly more efficient than previous NAS schemes which sample the architectures and filter them using the property predictors. We validate the effectiveness of DiffusionNAG through extensive experiments in two predictor-based NAS scenarios: Transferable NAS and Bayesian Optimization (BO)-based NAS. DiffusionNAG achieves superior performance with speedups of up to 20× when compared to the baselines on Transferable NAS benchmarks. Furthermore, when integrated into a BO-based algorithm, DiffusionNAG outperforms existing BO-based NAS approaches, particularly in the large MobileNetV3 search space on the ImageNet 1K dataset.
@@ 1 Introduction
While Neural Architecture Search (NAS) approaches automate neural architecture design, eliminating the need for manual design process with trial-and-error (Zoph & Le, 2017; Liu et al., 2019; Cai et al., 2019; Luo et al., 2018; Real et al., 2019; White et al., 2020), they mostly suffer from the high search cost, which often includes the full training with the searched architectures. To address this issue, many previous works have proposed to utilize parameterized property predictors (Luo et al., 2018; White et al., 2021a; b; 2023; Ning et al., 2020; 2021; Dudziak et al., 2020; Lee et al., 2021a; b) that can predict the performance of an architecture without training. However, existing NAS approaches still result in large waste of time as they need to explore an extensive search space and the property predictors mostly play a passive role such as the evaluators that rank architecture candidates provided by a search strategy to simply filter them out during the search process.
To overcome such limitations, we propose a paradigm shift from NAS (Neural Architecture Search) to a novel conditional NAG (Neural Architecture Generation) framework that enables the generation of desired neural architectures. Specifically, we introduce a novel predictor-guided Diffusion-based Neural Architecture Generative framework called DiffusionNAG, which explicitly incorporates the predictors into generating architectures that satisfy the objectives (e.g., high accuracy or robustness against attack). To achieve this goal, we employ the diffusion generative models (Ho et al., 2020; Song et al., 2021b), which generate data by gradually injecting noise into the data and learning to reverse this process. They have demonstrated remarkable generative performance across a wide range of domains. Especially, we are inspired by their parameterized model-guidance mechanism (Sohl-Dickstein et al., 2015; Vignac et al., 2022) that allows the diffusion generative models to excel in conditional generation over diverse domains such as generating images that match specific labels (Ramesh et al., 2021) or discovering new drugs meeting particular property criteria (Lee et al., 2023).
@@
We demonstrate the effectiveness of DiffusionNAG with extensive experiments under two key predictor-based NAS scenarios: 1) Transferable NAS and 2) Bayesian Optimization (BO)-based NAS. For Transferable NAS using transferable dataset-aware predictors, DiffusionNAG achieves superior or comparable performance with the speedup of up to 20$\times$ on four datasets from Transferable NAS benchmarks, including the large MobileNetV3 (MBv3) search space and NAS-Bench-201. Notably, DiffusionNAG demonstrates superior generation quality compared to MetaD2A (Lee et al., 2021a), a closely related unconditional generation-based method. For BO-based NAS with task-specific predictors, DiffusionNAG outperforms existing BO-based NAS approaches that rely on heuristic acquisition optimization strategies, such as random architecture sampling or architecture mutation, across four acquisition functions. This is because DiffusionNAG overcomes the limitation of existing BO-based NAS, which samples low-quality architectures during the initial phase, by sampling from the space of the architectures that satisfy the given properties. DiffusionNAG obtains especially large performance gains on the large MBv3 search space on the ImageNet 1K dataset, demonstrating its effectiveness in restricting the solution space when the search space is large. Furthermore, we verify that our score network generates 100% valid architectures by successfully capturing their computational flow, whereas the diffusion model for undirected graphs (Jo et al., 2022) almost fails.
@@ 2.1 Neural Architecture Diffusion Process
As a first step, we formulate an unconditional neural architecture diffusion process. Following Song et al. (2021b), we define a forward diffusion process that describes the perturbation of neural architecture distribution (search space) to a known prior distribution (e.g., Gaussian normal distribution) modeled by a stochastic differential equation (SDE), and then learn to reverse the perturbation process to sample the architectures from the search space starting from noise.
@@
The reverse-time diffusion process corresponding to the forward process is modeled by the following SDE (Anderson, 1982; Song et al., 2021b):
@@
In order to use the reverse process as a generative model, the score network $\bm{s}_{\bm{\theta}}$ is trained to approximate the score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})$ with the following score matching (Hyvärinen, 2005; Song et al., 2021a) objective, where $\lambda(t)$ is a given positive weighting function:
@@
where, $\mathbf{Emb}_{ops}\!\left({{\mathbf{v}}}_{i}\right)$ and $\mathbf{Emb}_{time}\!\left(t\right)$ are embeddings of each node (operation) ${{\mathbf{v}}}_{i}$ and time $t$, respectively.
@@ 2.2 Conditional Neural Architecture Generation
Inspired by the parameterized model-guidance scheme (Sohl-Dickstein et al., 2015; Vignac et al., 2022; Dhariwal & Nichol, 2021), we incorporate a parameterized predictor in our framework to actively guide the generation toward architectures that satisfy specific objectives. Let ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}$ be the desired property (e.g., high accuracy or robustness against attacks) we want the neural architectures to satisfy. Then, we include the information of ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}$ into the score function. To be specific, we generate neural architectures from the conditional distribution $p_{t}(\bm{A}_{t}|{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}})$ by solving the following conditional reverse-time SDE (Song et al., 2021b):
$\mathrm{d}\bm{A}_{t}=\left[\mathbf{f}_{t}(\bm{A}_{t})-g_{t}^{2}\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}})\right]\!\mathrm{d}\mkern 1.5mu\overline{\mkern-1.5mut\mkern-1.5mu}\mkern 1.5mu+g_{t}\mathrm{d}\bar{\mathbf{w}}.$ (6)
Here, we can decompose the conditional score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}})$ in Equation 6 as the sum of two gradients that is derived from the Bayes’ theorem $p(\bm{A}_{t}|{y})\propto p(\bm{A}_{t})\ p({y}|\bm{A}_{t})$:
$\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}})=\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})+\nabla_{\bm{A}_{t}}\!\log p_{t}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t}).$ (7)
By approximating the score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})$ with the score network $\bm{s}_{\theta^{*}}$, the conditional generative process of Equation 6 can be simulated if the term $\nabla_{\bm{A}_{t}}\!\log p_{t}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})$ could be estimated. Since $\log p_{t}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})$ represents the log-likelihood that the neural architecture $\bm{A}_{t}$ satisfies the target property ${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}$, we propose to model $\log p_{t}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})$ using a pre-trained predictor $f_{\bm{\phi}}(y|\bm{A}_{t})$ parameterized by $\bm{\phi}$, which predicts the desired property ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}$ given a perturbed neural architecture $\bm{A}_{t}$:
$\nabla_{\bm{A}_{t}}\!\log p_{t}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})\approx{\nabla_{\bm{A}_{t}}\!\log f_{\bm{\phi}}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})}.$ (8)
@@
$\mathrm{d}\bm{A}_{t}=\left\{\mathbf{f}_{t}(\bm{A}_{t})-g_{t}^{2}\Big{[}\bm{s}_{\bm{\theta}^{*}}(\bm{A}_{t},t)+k_{t}{\nabla_{\bm{A}_{t}}\!\log f_{\bm{\phi}}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y}}|\bm{A}_{t})}\Big{]}\right\}\mathrm{d}\mkern 1.5mu\overline{\mkern-1.5mut\mkern-1.5mu}\mkern 1.5mu+g_{t}\mathrm{d}\bar{\mathbf{w}}.$ (9)
@@ 3 Experiment
We validate the effectiveness of DiffusionNAG on two predictor-based NAS scenarios: Transferable NAS (Section 3.1) and BO-based NAS (Section 3.2). In Section 3.3, we demonstrate the effectiveness of the proposed score network. In Section 3.4, we analyze DiffusionNAG’s adaptation ability.
@@ 3.1 Comparison with Transferable NAS Methods
Transferable NAS methods (Shala et al., 2023; Lee et al., 2021a) are designed to leverage prior knowledge learned from previous NAS tasks, making NAS more practical on an unseen task. To achieve this, all Transferable NAS methods, including our DiffusionNAG, utilize a transferable dataset-aware accuracy predictor, as described in Section 2.3. The dataset-aware predictor is meta-trained on the meta-dataset provided by Lee et al. (2021a), which consists of $153,408/4,230$ meta-training tasks for MBv3/NB201, respectively. For more details, please refer to Lee et al. (2021a). MetaD2A (Lee et al., 2021a), which is the most closely related to our work, includes an unconditional architecture generative model that explicitly excludes the dataset-aware predictor during the generation process. Instead, MetaD2A needs to search for optimal architectures across multiple tasks, train these architectures to obtain their accuracy data, and use this costly accuracy collection to train its generative model. Besides, it uses the dataset-aware predictor only during the subsequent evaluation stage to rank the generated architectures. During the test phase, it first objective-unconditionally generates 500 candidate architectures and then evaluates the top architectures using its predictor. TNAS (Shala et al., 2023) enhances the meta-learned dataset-aware predictor’s adaptability to unseen datasets by utilizing BO with the deep-kernel GP strategy without involving any generation process (Please see Section C.2 for details of the baselines.). DiffusionNAG conditionally generates 500 architectures with the diffusion model guided by the dataset-aware predictor. Our generation process takes up to 4.35 GPU minutes on Tesla V100-SXM2. Finally, we select the top architectures sorted by the predictor among the generated candidates. We conduct experiments on Transferable NAS benchmarks (Lee et al., 2021a) such as four unseen datasets - CIFAR-10, CIFAR-100, Aircraft (Maji et al., 2013), and Oxford IIT Pets (Parkhi et al., 2012) from large search space MBv3 (Table 1) and, NB201 (Table 2).
@@
In  Table 1, MetaD2A, TNAS, and DiffusionNAG obtain the top 30 neural architectures for each datasets, following the descriptions in the Experimental Setup section. Subsequently, we train these architectures on the datasets following the training pipeline described in Section C.5. Once the architectures are trained, we analyze the accuracy statistics for each method’s group of architectures. Additionally, we calculate p-value to assess the statistical significance of performance differences between the architecture groups obtained via DiffusionNAG and each method. A p-value of 0.05 or lower denotes that a statistically meaningful difference exists in the performances of the generated architectures between the two groups.
@@
We highlight two key aspects from the results of Table 2. Firstly, the architectures generated by DiffusionNAG attain oracle accuracies of $\mathbf{94.37\%}$ and $\mathbf{73.51\%}$ on CIFAR-10 and CIFAR-100 datasets, respectively, and outperform architectures obtained by the baseline methods on Aircraft and Oxford-IIIT Pets datasets. While MetaD2A and TNAS achieve accuracies of $\mathbf{59.15\%/57.71\%}$ and $\mathbf{40.00\%/39.04\%}$ on Aircraft and Oxford-IIIT Pets datasets, respectively, DiffusionNAG gains even better accuracies of $\mathbf{59.63\%}$ and $\mathbf{41.32\%}$, demonstrating its superiority. Secondly, DiffusionNAG significantly improves the search efficiency by minimizing the number of architectures that require full training (Trained Archs) to obtain a final accuracy (For CIFAR-10 and CIFAR-100, an accuracy is retrieved from NB201 benchmarks) compared to all baselines. Specifically, when considering the Aircraft and Oxford-IIIT Pets datasets, DiffusionNAG only needs to train $\mathbf{2}$ architectures for each dataset to complete the search process while MetaD2A and TNAS require $\mathbf{40/26}$ and $\mathbf{6/40}$ architectures, respectively. This results in a remarkable speedup of at least $\mathbf{8.4\times}$ and up to $\mathbf{20\times}$ on average.

### 24 Mar 2024 v4

@@ Abstract
Existing NAS methods suffer from either an excessive amount of time for repetitive sampling and training of many task-irrelevant architectures. To tackle such limitations of existing NAS methods, we propose a paradigm shift from NAS to a novel conditional Neural Architecture Generation (NAG) framework based on diffusion models, dubbed DiffusionNAG. Specifically, we consider the neural architectures as directed graphs and propose a graph diffusion model for generating them. Moreover, with the guidance of parameterized predictors, DiffusionNAG can flexibly generate task-optimal architectures with the desired properties for diverse tasks, by sampling from a region that is more likely to satisfy the properties. This conditional NAG scheme is significantly more efficient than previous NAS schemes which sample the architectures and filter them using the property predictors. We validate the effectiveness of DiffusionNAG through extensive experiments in two predictor-based NAS scenarios: Transferable NAS and Bayesian Optimization (BO)-based NAS. DiffusionNAG achieves superior performance with speedups of up to 35$\times$ when compared to the baselines on Transferable NAS benchmarks. Furthermore, when integrated into a BO-based algorithm, DiffusionNAG outperforms existing BO-based NAS approaches, particularly in the large MobileNetV3 search space on the ImageNet 1K dataset. Code is available at https://github.com/CownowAn/DiffusionNAG.
@@ 1 Introduction
While Neural Architecture Search (NAS) approaches automate neural architecture design, eliminating the need for manual design process with trial-and-error (Zoph & Le, 2017; Liu et al., 2019; Cai et al., 2019; Luo et al., 2018; Real et al., 2019; White et al., 2020), they mostly suffer from the high search cost, which often includes the full training with the searched architectures. To address this issue, many previous works have proposed to utilize parameterized property predictors (Luo et al., 2018; White et al., 2021a; b; 2023; Ning et al., 2020; 2021; Dudziak et al., 2020; Lee et al., 2021a; b; 2023a) that can predict the performance of an architecture without training. However, existing NAS approaches still result in large waste of time as they need to explore an extensive search space and the property predictors mostly play a passive role such as the evaluators that rank architecture candidates provided by a search strategy to simply filter them out during the search process.
To overcome such limitations, we propose a paradigm shift from NAS (Neural Architecture Search) to a novel conditional NAG (Neural Architecture Generation) framework that enables the generation of desired neural architectures. Specifically, we introduce a novel predictor-guided Diffusion-based Neural Architecture Generative framework called DiffusionNAG, which explicitly incorporates the predictors into generating architectures that satisfy the objectives (e.g., high accuracy or robustness against attack). To achieve this goal, we employ the diffusion generative models (Ho et al., 2020; Song et al., 2021), which generate data by gradually injecting noise into the data and learning to reverse this process. They have demonstrated remarkable generative performance across a wide range of domains. Especially, we are inspired by their parameterized model-guidance mechanism (Sohl-Dickstein et al., 2015; Vignac et al., 2022) that allows the diffusion generative models to excel in conditional generation over diverse domains such as generating images that match specific labels (Ramesh et al., 2021) or discovering new drugs meeting particular property criteria (Lee et al., 2023b).
@@
We demonstrate the effectiveness of DiffusionNAG with extensive experiments under two key predictor-based NAS scenarios: 1) Transferable NAS and 2) Bayesian Optimization (BO)-based NAS. For Transferable NAS using transferable dataset-aware predictors, DiffusionNAG achieves superior or comparable performance with the speedup of up to 35$\times$ on four datasets from Transferable NAS benchmarks, including the large MobileNetV3 (MBv3) search space and NAS-Bench-201. Notably, DiffusionNAG demonstrates superior generation quality compared to MetaD2A (Lee et al., 2021a), a closely related unconditional generation-based method. For BO-based NAS with task-specific predictors, DiffusionNAG outperforms existing BO-based NAS approaches that rely on heuristic acquisition optimization strategies, such as random architecture sampling or architecture mutation, across four acquisition functions. This is because DiffusionNAG overcomes the limitation of existing BO-based NAS, which samples low-quality architectures during the initial phase, by sampling from the space of the architectures that satisfy the given properties. DiffusionNAG obtains especially large performance gains on the large MBv3 search space on the ImageNet 1K dataset, demonstrating its effectiveness in restricting the solution space when the search space is large. Furthermore, we verify that our score network generates 100% valid architectures by successfully capturing their computational flow, whereas the diffusion model for undirected graphs (Jo et al., 2022) almost fails.
@@ 2.1 Neural Architecture Diffusion Process
As a first step, we formulate an unconditional neural architecture diffusion process. Following Song et al. (2021), we define a forward diffusion process that describes the perturbation of neural architecture distribution (search space) to a known prior distribution (e.g., Gaussian normal distribution) modeled by a stochastic differential equation (SDE), and then learn to reverse the perturbation process to sample the architectures from the search space starting from noise.
@@
The reverse-time diffusion process corresponding to the forward process is modeled by the following SDE (Anderson, 1982; Song et al., 2021):
@@
In order to use the reverse process as a generative model, the score network $\bm{s}_{\bm{\theta}}$ is trained to approximate the score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})$ with the following score matching (Hyvärinen, 2005; Song et al., 2021) objective, where $\lambda(t)$ is a given positive weighting function:
@@
where $\mathbf{Emb}_{ops}\!\left({{\mathbf{v}}}_{i}\right)$ and $\mathbf{Emb}_{time}\!\left(t\right)$ are embeddings of each node ${{\mathbf{v}}}_{i}$ and time $t$, respectively.
@@ 2.2 Conditional Neural Architecture Generation
Inspired by the parameterized model-guidance scheme (Sohl-Dickstein et al., 2015; Vignac et al., 2022; Dhariwal & Nichol, 2021), we incorporate a parameterized predictor in our framework to actively guide the generation toward architectures that satisfy specific objectives. Let $y$ be the desired property (e.g., high accuracy or robustness against attacks) we want the neural architectures to satisfy. Then, we include the information of $y$ into the score function. To be specific, we generate neural architectures from the conditional distribution $p_{t}(\bm{A}_{t}|{y})$ by solving the following conditional reverse-time SDE (Song et al., 2021):
$\mathrm{d}\bm{A}_{t}=\left[\mathbf{f}_{t}(\bm{A}_{t})-g_{t}^{2}\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{y})\right]\!\mathrm{d}\mkern 1.5mu\overline{\mkern-1.5mut\mkern-1.5mu}\mkern 1.5mu+g_{t}\mathrm{d}\bar{\mathbf{w}}.$ (6)
Here, we can decompose the conditional score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{y})$ in Equation 6 as the sum of two gradients that is derived from the Bayes’ theorem $p(\bm{A}_{t}|{y})\propto p(\bm{A}_{t})\ p({y}|\bm{A}_{t})$:
$\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t}|{y})=\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})+\nabla_{\bm{A}_{t}}\!\log p_{t}({y}|\bm{A}_{t}).$ (7)
By approximating the score function $\nabla_{\bm{A}_{t}}\!\log p_{t}(\bm{A}_{t})$ with the score network $\bm{s}_{\theta^{*}}$, the conditional generative process of Equation 6 can be simulated if the term $\nabla_{\bm{A}_{t}}\!\log p_{t}({y}|\bm{A}_{t})$ could be estimated. Since $\log p_{t}({y}|\bm{A}_{t})$ represents the log-likelihood that the neural architecture $\bm{A}_{t}$ satisfies the target property ${y}$, we propose to model $\log p_{t}({y}|\bm{A}_{t})$ using a pre-trained predictor $f_{\bm{\phi}}(y|\bm{A}_{t})$ parameterized by $\bm{\phi}$, which predicts the desired property $y$ given a perturbed neural architecture $\bm{A}_{t}$:
$\nabla_{\bm{A}_{t}}\!\log p_{t}({y}|\bm{A}_{t})\approx{\nabla_{\bm{A}_{t}}\!\log f_{\bm{\phi}}({y}|\bm{A}_{t})}.$ (8)
@@
$\mathrm{d}\bm{A}_{t}=\left\{\mathbf{f}_{t}(\bm{A}_{t})-g_{t}^{2}\Big{[}\bm{s}_{\bm{\theta}^{*}}(\bm{A}_{t},t)+k_{t}{\nabla_{\bm{A}_{t}}\!\log f_{\bm{\phi}}({y}|\bm{A}_{t})}\Big{]}\right\}\mathrm{d}\mkern 1.5mu\overline{\mkern-1.5mut\mkern-1.5mu}\mkern 1.5mu+g_{t}\mathrm{d}\bar{\mathbf{w}}.$ (9)
@@ 3 Experiment
We validate the effectiveness of DiffusionNAG on two predictor-based NAS scenarios: Transferable NAS (Section 3.1) and BO-based NAS (Section 3.2). In Section 3.3, we demonstrate the effectiveness of the proposed score network.
@@ 3.1 Comparison with Transferable NAS Methods
Transferable NAS methods (Shala et al., 2023; Lee et al., 2021a) are designed to leverage prior knowledge learned from previous NAS tasks, making NAS more practical on an unseen task. To achieve this, all Transferable NAS methods, including our DiffusionNAG, utilize a transferable dataset-aware accuracy predictor, as described in Section 2.3. The dataset-aware predictor is meta-trained on the meta-dataset provided by Lee et al. (2021a), which consists of $153,408/4,230$ meta-training tasks for MBv3/NB201, respectively. For more details, please refer to Lee et al. (2021a). MetaD2A (Lee et al., 2021a), which is the most closely related to our work, includes an unconditional architecture generative model that explicitly excludes the dataset-aware predictor during the generation process. Instead, MetaD2A needs to search for optimal architectures across multiple tasks, train these architectures to obtain their accuracy data, and use this costly accuracy collection to train its generative model. Besides, it uses the dataset-aware predictor only during the subsequent evaluation stage to rank the generated architectures. During the test phase, it first objective-unconditionally generates architectures and then evaluates the top architectures using its predictor. TNAS (Shala et al., 2023) enhances the meta-learned dataset-aware predictor’s adaptability to unseen datasets by utilizing BO with the deep-kernel GP strategy without involving any generation process (Please see Section C.2 for details of the baselines.). DiffusionNAG conditionally generates architectures with the diffusion model guided by the dataset-aware predictor. Our generation process, with a sampling batch size of 256, takes up to 2.02 GPU minutes on Tesla V100-SXM2 to sample one batch. Finally, we select the top architectures sorted by the predictor among the generated candidates. We conduct experiments on Transferable NAS benchmarks (Lee et al., 2021a) such as four unseen datasets - CIFAR-10, CIFAR-100, Aircraft (Maji et al., 2013), and Oxford IIT Pets (Parkhi et al., 2012) from large search space MBv3 (Table 1) and, NB201 (Table 2).
@@
In  Table 1, MetaD2A, TNAS, and DiffusionNAG obtain the top 30 neural architectures for each datasets. Subsequently, we train these architectures on the datasets following the training pipeline described in Section C.5. Once the architectures are trained, we analyze the accuracy statistics for each method’s group of architectures. Additionally, we calculate p-value to assess the statistical significance of performance differences between the architecture groups obtained via DiffusionNAG and each method. A p-value of 0.05 or lower denotes that a statistically meaningful difference exists in the performances of the generated architectures between the two groups.
@@
[…]